OpenStudy (anonymous):

determine the angle through which the minute hand of a clock travels between 9:18am and 11:55am on the same morning?

OpenStudy (saifoo.khan):

1 min = 360/60 = 6degrees. total minutes, 60+60+37 = 157 therefore, 157*6 = 942degrees/

OpenStudy (anonymous):

thanx

OpenStudy (saifoo.khan):

np